OpenStudy (newwar):

What is the value of f(−1)f(−1) when f(x)=2x+2f(x)=2x+2 ? Enter your answer in the box. f(−1)=f(−1)=

jhonyy9 (jhonyy9):

there is two time f(-1) or what mean this ?

jhonyy9 (jhonyy9):

and f(x) same too ,why ?

jhonyy9 (jhonyy9):

so than there is f(x) = 2x+2 and you need calculi the f(-1) = ? so just substitute in place of x the -1 and make the calculi so will get the necessary right answer

jhonyy9 (jhonyy9):

ok. ?

OpenStudy (newwar):

ok

jhonyy9 (jhonyy9):

good luck was my pleasure bye bye

OpenStudy (newwar):

ok i did it was 7 right or did i do it wrong

jhonyy9 (jhonyy9):

for you get f(-1) just substitute in place of x the -1 f(-1) = 2(-1) +2 right ?

jhonyy9 (jhonyy9):

write me please how you ve got this 7 ?

OpenStudy (newwar):

o i am sorry i got 0 @jhonyy9

jhonyy9 (jhonyy9):

yes bc. 2*(-1) = -2 +2 = 0 hope helped

OpenStudy (newwar):

yes thanks

jhonyy9 (jhonyy9):

yw my pleasure good luck